40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14502     Entry time: Fri Mar 29 21:00:06 2019     In reply to: 14498     Reply to this: 14503
Author: gautam 
Type: Update 
Category: ALS 
Subject: BeatMouth with NF1611s installed 
  • Newfocus 15V current limited supply was taken from bottom NE corner of the ITMY Oplev table to power the BeatMouth on the PSL table
  • BeatMouth was installed on top shelf on PSL table [Attachment #1].
  • Light levels in fibers were checked:
    • PSL: initially, only ~200uW / 4mW was coupled in. This was improved to 2.6mW/4mW (~65% MM) which was deemed sufficient for a first test), by tweaking the alignment of, and into the collimator.
    • EX: ~900uW measured at the PSL table. I remember the incident power being ~1mW. So this is pretty good.
  • Fibers hooked up to BeatMouth:
    • EX light only, DC mon of X PD reads -2.1V.
    • With PSL light, I get -4.6 V.
    • For these numbers, with the DC transimpedance of 10kohm and the RF transimpedance of 700 ohm, I expect a beat of ~0dBm
  • DC light level stability is being monitored by a temporarily hijacked PSL NPRO diagnostic Acromag channel. Main motivation is to confirm that the alignment to the special axes of the PM fibers is still good and we aren't seeing large tempreature-driven waveplate effects.
  • RF part of the circuit is terminated into 50ohms for now -
    • there is still a quesiton as to what is the correct RF amplifier to use in sending the signal to the 1Y3 rack.
    • An initial RF beat power level measurement yielded -5dBm, which is inconsistent with the DC monitor voltages, but I'm not sure what frequency the beat was at, will make a more careful measurement with a scope or the network analyzer.
    • We want the RF pre-amp to be:
      • Low noise, keeping this in mind
      • High enough gain to boost the V/Hz discriminant of the electronic delay line
      • Not too high gain that we run into compression / saturate some of the delay line electronics - specifically, the LO input of the LSC demod board has a Teledyne amp in the signal chain, and so we need to ensure the signal level there is <16dBm (nominal level is 10dBm).
      • I'm evaluating options...
  • At 1Y3:
    • I pulled out the delay-line enclosure, and removed the (superglued) resistive power splitters with the help of some acetone
    • The newly acquired power splitters (ZAPD-2-252-S+) were affixed to the front panel, in which I made some mounting holes.
    • The new look setup, re-installed at 1Y3, is shown in Attachment #2.
Attachment 1: IMG_7384.JPG  1.198 MB  Uploaded Sun Mar 31 15:58:51 2019  | Hide | Hide all
IMG_7384.JPG
Attachment 2: IMG_7385.JPG  1.358 MB  Uploaded Sun Mar 31 16:02:44 2019  | Hide | Hide all
IMG_7385.JPG
ELOG V3.1.3-